# RoTipBot: Robotic Handling of Thin and Flexible Objects using Rotatable Tactile Sensors

### v1

@@ Abstract
This paper introduces RoTipBot, a novel robotic system for handling thin, flexible objects. Different from previous works that are limited to singulating them using suction cups or soft grippers, RoTipBot can grasp and count multiple layers simultaneously, emulating human handling in various environments. Specifically, we develop a novel vision-based tactile sensor named RoTip that can rotate and sense contact information around its tip. Equipped with two RoTip sensors, RoTipBot feeds multiple layers of thin, flexible objects into the centre between its fingers, enabling effective grasping and counting. RoTip’s tactile sensing ensures both fingers maintain good contact with the object, and an adjustment approach is designed to allow the gripper to adapt to changes in the object. Extensive experiments demonstrate the efficacy of the RoTip sensor and the RoTipBot approach. The results show that RoTipBot not only achieves a higher success rate but also grasps and counts multiple layers simultaneously – capabilities not possible with previous methods. Furthermore, RoTipBot operates up to three times faster than state-of-the-art methods. The success of RoTipBot paves the way for future research in object manipulation using mobilised tactile sensors. All the materials used in this paper are available at https://sites.google.com/view/rotipbot.
@@ I Introduction
Thin, flexible objects are common in our daily lives, e.g., plastic wrap in kitchens, printer paper in offices and delicate lab gloves. Handling these objects is becoming increasingly important for assistive robots, driven by the growing demand for their use in household and everyday settings. For example, assistive robots need to be able to turn pages with precision to help the elderly read and use plastic wrap to store food. Therefore, the development of robotic systems capable of handling these thin, flexible materials with precision is essential for the next generation of assistive robots.
@@
However, thin and flexible objects, characterised by their relatively small thickness compared to their length and width, and their ability to bend or flex without breaking, present two primary challenges for assistive robots. First, their thinness often results in layers that stack over each other, making the states beneath the top layer difficult to discern with a camera, resulting into incomplete and noisy perception [1, 2]. On the other hand, the tendency of such objects to deform may require dexterity and compliance in robotic grippers [3, 2, 4].
@@
Figure 1: (a) A demonstration of RoTipBot. The tactile sensors ensure good contact with objects, while the rotation capability feeds multiple layers of thin, flexible objects into the centre for grasping and counting. Different transparencies of the paper represent states at different time steps. (b) Snapshots of the feeding process for multiple print papers. (c-e) Sketches comparing RoTipBot to approaches based on suction cups and soft grippers. RoTipBot can grasp and count multiple layers, whereas the other methods cannot.
@@
Existing approaches for grasping thin, flexible objects often bypass these two challenges. Methods typically assume objects are in known positions [5], or rely on costly force/torque sensors to compensate for the noise in visual perception [1], significantly increasing the overall cost. For the second challenge, specialised mechanisms like suction [6, 7] and two-finger soft grippers [8, 4, 1] are used. However, these mechanisms allow only one page to be grasped at a time, limiting efficiency. Suction cups struggle with air gaps between sheets, preventing effective vacuum seals, while two-finger soft grippers [8, 4, 1] lack the dexterity for continuous picking and holding multiple sheets securely.
@@
To address the above challenges in handling thin and flexible objects, in this paper we propose a novel robotic approach named RoTipBot that uses rotatable tactile sensors to sense the contacts with these objects and to feed multiple layers of these objects into the centre for grasping and counting. We develop a new vision-based tactile sensor named RoTip, which can sense the entire fingertip area and actively rotate its body. As shown in Fig. 1-(a), RoTip’s sensing capability mitigates visual perception noise and ensures good contact with objects. Its rotational capability adds an extra Degree of Freedom, enhancing dexterity in handling thin and flexible objects. A segmentation and projection model is developed to obtain contact areas and surface planes to guide grasping. As shown in Fig. 1-(b), the passive finger of RoTipBot holds the object while the active finger rotates to feed multiple layers into the centre. To adapt to the changes in stacked thickness, the fingers continuously adjust their pose to maintain contact and ensure accurate control when grasping multiple layers.
@@
  • We introduce RoTipBot, a novel robot for handling thin and flexible objects that can grasp and count them for the first time.
  • The RoTip sensors in RoTipBot can sense contacts across the entire fingertip surface, with an additional motor enhancing dexterity by actively rotating the sensor to slide along the object’s surface.
  • We conducted comprehensive experiments to evaluate RoTipBot, demonstrating its ability to grasp and count multiple layers of thin and flexible objects, with high success rates and efficiency.
@@ II-A Vision-based Tactile Sensors
In recent years, many vision-based tactile sensors have been developed, primarily classified into two main categories: the GelSight family [9], which processes raw images and employs photometric stereo for shape reconstruction; the TacTip family [10], which estimates contact by tracking pin displacement. Unlike TacTip, which is limited by pin resolution, GelSight offers superior resolution and detailed texture information [11]. Our RoTip sensor belongs to the GelSight family.
The first GelSight prototype was developed in 2009 by Johnson and Adelson [12]. Following that, a set of subsequent developments were built with different focuses, such as enhanced lighting setups of GelSight-2017 [9] and GelSlim 3.0 [13], sensor reproducibility of DIGIT [14], and a compact shape for robot fingers of [15]. While previous GelSight tactile sensors capture high-resolution tactile images, their limited sensing area poses challenges for complex robotic manipulation. To address this, recent works have introduced finger-shaped tactile sensors, such as GelTip [16], OmniTact [17], Insight [18] and GelSight360 [19]. These sensors can be mounted on off-the-shelf parallel grippers to enhance manipulation tasks with tactile sensing.
However, these setups possess limited degrees of freedom, decoupling tactile sensing from robotic dexterity. Hence, several studies have explored the integration of rotational movement into tactile sensors to enhance functionality. TouchRoller [20] was proposed to rapidly reconstruct large contact surfaces through surface rotation, while TacRot [21] utilises a rotatable surface to facilitate in-hand object manipulation. The tactile-reactive roller grasper (TRRG) [22] integrates compliant, steerable cylindrical fingertips for improved manipulation. Despite these advancements, they only perceive the inner side of the fingertip, preventing interaction modes such as pressing or poking. To our best knowledge, RoTip is the first tactile sensor that can sense the entire fingertip area and actively rotate its finger body. These two capabilities are crucial for addressing the challenges of incomplete visual perception and the need for dexterity in handling thin and flexible objects.
@@ II-B Thin and Flexible Object Grasping
Deformable object manipulation has been a growing research area with various applications, such as assembling cables in factories [23, 24, 25, 26], assisted dressing and laundry folding [27, 28, 29, 30] and food production [31, 32]. These tasks mainly rely on visual perception [23, 33, 28, 32, 34]. However, the specific challenge of manipulating thin and flexible objects remains relatively unexplored. The small thickness of these objects compared to their length and width and their stacked layers complicates visual perception and grasping.
Recent approaches to manipulating thin and flexible objects focus on the control design. Several studies achieve thin object grasping via pre-grasp motions such as sliding [35] and scooping [36, 37]. In [35], thin objects are slid to the table edge and grasped from the side. In [36], stable grasps of thin objects are achieved through scooping actions. In [37], a prying strategy is proposed to tilt a thin object like a card to create clearance beneath it for a stable pinch grasp. Similarly, in [38], an underactuated crawler is used to tilt thin objects for pinch grasps.
@@
Another mechanism used for grasping delicate materials is the soft gripper. As discussed in [4], the three-axis compliance of a soft gripper provides unique benefits for grasping thin, flexible objects. In [8], a two-finger soft gripper is used to deform an object for a pinch grasp, but this approach requires knowledge of physical properties, such as friction, to model object deformation dynamics. Building on this, FlipBot introduced in [1] uses a two-finger soft gripper with a force/torque sensor to singulate and grasp thin and flexible objects by integrating exteroceptive and proprioceptive perceptions. However, it uses a force/torque sensor that only provides contact information at a single point, and can only grasp one page at a time, limiting its efficiency. In contrast, our RoTipBot offers rich distributed contact information using vision-based tactile sensing, and its rotation capability enables the grasping of multiple thin, flexible objects simultaneously. Another related work, [5], leverages tactile feedback to handle thin and flexible objects like paper, but simplifies the problem by placing paper on the edge of the table for easy pickup, bypassing the challenge of grasping from the table surface. Our RoTipBot addresses these issues effectively, proving its capability in these complex tasks.
III The RoTip sensor
In this section, we introduce RoTip, a novel vision-based tactile sensor that can sense the entire fingertip area and rotate along the vertical axis, as shown in Fig. 2. The sensor can be easily attached to different commercial grippers, such as the Robotiq-2f Gripper, as shown in Fig. 1-(a). We provide detailed descriptions of both the hardware and firmware to enable the research community to reproduce the sensor. A GitHub repository containing all necessary files has been prepared for interested readers.
III-A Sensor Design
@@
The proposed RoTip sensor comprises three primary modules: the fixing module, the transmission module and the finger body, as shown in Fig. 2-(b), (c) and (d), respectively. The fixing module is used for secure attachment and precise positioning of the transmission module and the finger body, ensuring stability during operation. Meanwhile, the transmission module is used to drive the rotational motion, enabling dynamic functionality. Lastly, the finger body serves as the tactile interface, providing the sensor’s tactile perception capabilities. Further insights into these three modules are elaborated below:
III-A1 Fixing module
The fixing module of the RoTip sensor is designed for connecting and bolstering other components of the sensor, as shown in Fig 2-(b). It features a gripper connector, allowing for easy and secure attachment to a diverse range of robotic grippers. Both threaded mechanisms and screws are used to link the gripper connector and upper shell, so as to support forces and torques during interactions with objects. Moreover, the module incorporates a dedicated support structure tailored for the Printed Circuit Board (PCB) of a camera, ensuring the camera’s stability. Additionally, the bottom section of the fixing module is designed for camera connection, ensuring that the camera maintains its position even when the finger body rotates. To protect the camera’s functionality, the camera cable is neatly channelled through a tube within the fixing module.
III-A2 Transmission module
The transmission module in our RoTip sensor features a gear transmission mechanism as shown in Fig. 2-(c). The Degree of Freedom for each finger is managed by a Robotis Dynamixel XM430-W350 actuator [^1] https://robotis.co.uk/dynamixel-xm430-w350-r.html, capable of executing different control modes: precise position control mode from 0 to 360 and velocity control mode for endless turns. To optimise performance, a transmission ratio of 5:3 was chosen to balance torque and speed, ensuring both efficient and responsive movements. This ratio not only enhances operational smoothness but also mitigates gear collision risks when the gripper is closed. Moreover, two 6803-2RS ball bearings, each sized at $17mm\times 26mm\times 5mm$, are utilised in the design to facilitate smooth rotational movement and provide stability to the moving parts of the system. For software integration, we utilise the DynamixelSDK, which is programmed in Python. This choice facilitates seamless integration with various deep-learning approaches, enhancing the sensor’s overall adaptability and versatility.
Figure 3: (a) and (b) represent RoTip sensor with a hemisphere tip and a flat tip, respectively. In the first two rows, the transparent support shells, the coated elastomer, and the corresponding tactile images are presented from left to right. Notably, sparking points in the centre of the tactile image are discernible due to the intense reflection from the flat tip. (c) showcases the tactile readings captured as the RoTip sensor with a hemisphere tip interacts with a bolt, a nut, and a USB cable.
III-A3 Finger body
The finger body serves as a tactile interface, resembling a fingertip, crucial for interacting with the environment. Following the design principles of our previous vision-based tactile sensor GelTip [16], our RoTip incorporates an elastomer with an opaque coating layer to enhance tactile feedback, along with a rigid transparent shell that provides structural support to the elastomer. To accommodate the camera, we set the radius of the rigid shell at 28 mm with a thickness of 2 mm to ensure sufficient hardness. Threaded mounting was employed to not only enhance the sensor’s stability but also facilitate replicability. This modular design allows for easy interchangeability of finger bodies, including alternative shapes like the flat-tipped variant shown in Fig. 3-(b) instead of the hemisphere tip.
To achieve uniform illumination, we designed a PCB with multiple LEDs arranged in a circular pattern as shown in Fig. 2-(d). Following [18], we incorporated a collimator over the LED PCB. Additionally, a diffuser was used to mitigate lighting sparks and enhance contrast and uniformity of tactile imprints, as suggested in [15]. The combination of photometric effects and structured lighting empowers the camera to precisely detect the deformations of the elastomer when it comes into physical contact with objects, such as the thread of a bolt, as illustrated in Fig. 3-(c).
Remark: For effective grasping of thin and flexible objects, it is essential that the RoTip sensors mounted on the gripper make contact during closure. To this end, the finger radius must be larger than that of any other component, including the transmission device and the shell.
Figure 4: Fabrication of RoTip’s fingertip. There are three main steps for this fabrication process: (a) 3D printing of moulds using the FormLab Form 3 3D printer; (b) Silicone fabrication through mixing XP-565 Silicone with the slacker to obtain the clear elastomer; (c) Coated layer fabrication through mixing XP565 with NOVOCS solvent, aluminium powder, and Silver Bullet powder; and finally (d) Assembled fingertip.
III-B Sensor Fabrication
In the fabrication of RoTip sensors, several components, such as the camera and LED PCB, are commercially available. For instance, we utilise an OV5640 camera [^2] https://www.waveshare.com/ov5640-camera-board-a.htm with a 160-degree field of view to capture tactile images. This camera is fitted into a compact camera holder, which allows for stable placement and orientation adjustment. The LED PCB is sourced from a specialised manufacturer, and detailed specifications are available on our website. However, for other critical parts such as the finger body, we have to manually design and create custom moulds and fabricate its rigid shell and elastomer.
III-B1 Transparent fingertip shell
We adopt a fabrication process utilising Stereolithography (SLA) 3D printing, in contrast to our previous approach in [16] that used a standard transparent test tube. This new method involves creating the shell with clear resin, which is subsequently refined through surface polishing and lacquer coating [39], resulting in a remarkably transparent shell as shown in Fig. 3. Additionally, we have adopted threading for assembling the fingertip shell, which enhances stability and eliminates the complexities associated with using screws. It is important to note that applying a lacquer coating slightly increases the thickness of the thread. As a result, the component designated for mounting the rigid shell needs to be adjusted to accommodate this slight change in dimension.
III-B2 Elastomer
We use the same materials in our previous work [16], i.e., XP-565 from Silicones, Inc. (High Point, NC, USA) and Slacker from Smooth-on Company. After extensive experiments we found the best ratios to be 1:15:15, i.e., 1 gram of XP-565 part A, 15 grams of XP-565 part B and 15 grams of the Slacker. This amount of mixture is sufficient to fabricate four sensor membranes. The elastomer is fabricated through moulding with a two-piece mould, as shown in Fig. 4. To fabricate the rigid parts, we take advantage of 3D printing technology, specifically SLA printing, due to its ability to produce much smoother mould surfaces, significantly impacts the quality of the elastomer. A thorough polishing step is also necessary to eliminate the layer lines from the 3D printing process.
The exterior of the elastomer is coated with an opaque reflective material aimed to contain and intensify the lighting within the shell. Aluminium powder with reflective pigment was used for coatings, as suggested in [9]. Specifically, the optimal ratio for the coating mixture is 1:15:4:4:10, i.e., 1 gram of XP-565 part A, 15 grams of XP-565 part B, 4 grams of aluminium powder, 4 grams of Silver Cast Magic® from the Smooth-on Company, and 10 grams of NOVOCS Matte Solvent. Then the cured silicone base was dipped into this mixture to form the opaque reflective coating layer. The acquired coating has proven to be durable and reliable under high contact forces. Figure 3-(c) demonstrates high-resolution and clear tactile images captured during contact with various objects, e.g., a bolt, a nut and a USB cable.

### v2

@@ Abstract
This paper introduces RoTipBot, a novel robotic system for handling thin, flexible objects. Different from previous works that are limited to singulating them using suction cups or soft grippers, RoTipBot can count multiple layers and then grasp them simultaneously in a single grasp closure. Specifically, we first develop a vision-based tactile sensor named RoTip that can rotate and sense contact information around its tip. Equipped with two RoTip sensors, RoTipBot rolls and feeds multiple layers of thin, flexible objects into the centre between its fingers, enabling effective grasping. Moreover, we design a tactile-based grasping strategy that uses RoTip’s sensing ability to ensure both fingers maintain secure contact with the object while accurately counting the number of fed objects. Extensive experiments demonstrate the efficacy of the RoTip sensor and the RoTipBot approach. The results show that RoTipBot not only achieves a higher success rate but also grasps and counts multiple layers simultaneously – capabilities not possible with previous methods. Furthermore, RoTipBot operates up to three times faster than state-of-the-art methods. The success of RoTipBot paves the way for future research in object manipulation using mobilised tactile sensors. All the materials used in this paper are available at https://sites.google.com/view/rotipbot.
@@ I Introduction
Thin, flexible objects are common in our daily lives, e.g., plastic wraps in kitchens, printer papers in offices and delicate lab gloves. Handling these objects is becoming increasingly important for assistive robots, driven by the growing demand for their use in household and everyday settings. For example, assistive robots need to be able to turn pages with precision to help the elderly read and use plastic wraps to store food. Therefore, the development of robotic systems capable of handling these thin, flexible materials with precision is essential for the next generation of assistive robots.
@@
However, thin and flexible objects, characterised by their relatively small thickness compared to their length and width, and their ability to bend or flex without breaking, present two primary challenges for assistive robots. First, their thinness often results in layers that stack over each other, making the states beneath the top layer difficult to discern with a camera, resulting in incomplete and noisy perception [1, 2]. On the other hand, the tendency of such objects to deform may require dexterity and compliance in robotic grippers [3, 2, 4].
@@
Figure 1: (a) A demonstration of RoTipBot. The tactile sensors ensure good contact with objects, while the rotation capability feeds multiple layers of thin, flexible objects into the centre for grasping and counting. Different transparencies of the paper represent states at different time steps. (b) Snapshots of the feeding process for multiple print papers. (c-e) Sketches comparing RoTipBot to approaches based on suction cups and soft grippers. RoTipBot can count multiple layers and then grasp them simultaneously in a single grasp closure, whereas the other methods cannot.
@@
Existing approaches for grasping thin, flexible objects often bypass these two challenges. Methods typically assume objects are in known positions [5], or rely on costly force/torque sensors to compensate for the noise in visual perception [1], significantly increasing the overall cost. For the second challenge, specialised mechanisms like suction [6, 7] and two-finger soft grippers [8, 4, 1] are used. However, these mechanisms allow only one page in each grasp closure, limiting efficiency. Suction cups struggle with air gaps between sheets, preventing effective vacuum seals, while two-finger soft grippers [8, 4, 1] lack the dexterity for multi-page grasping where the system first counts multiple layers and then grasps them simultaneously in a single grasp closure.
@@
To address the above challenges in handling thin and flexible objects, in this paper we propose a novel robotic approach named RoTipBot that uses rotatable tactile sensors to sense the contacts with these objects and to feed multiple layers of these objects into the centre for grasping and counting. We develop a new vision-based tactile sensor named RoTip, which can sense the entire fingertip area and actively rotate its body. As shown in Fig. 1-(a), RoTip’s sensing capability mitigates visual perception noise and ensures good contact with objects. Its rotational capability adds an extra Degree of Freedom, enhancing dexterity in handling thin and flexible objects. A segmentation and projection model is developed to obtain contact areas and surface planes to guide grasping. As shown in Fig. 1-(b), the passive finger of RoTipBot holds the object while the active finger rotates to feed multiple layers into the centre. To adapt to the changes in stacked thickness, the fingers continuously adjust their pose to maintain contact and ensure accurate control when grasping multiple layers. Additionally, tactile sensing is employed for accurate counting of the number of grasped papers.
@@
  • We introduce RoTipBot, a novel approach to handling multiple thin, flexible objects, integrating tactile sensing with high dexterity and accuracy.
  • We design RoTip sensors that can sense contacts across the entire fingertip surface with an additional motor enabling active rotation of the sensor surface without angle constraints, significantly enhancing the system’s flexibility and precision in object manipulation.
  • We propose a tactile-based grasping strategy that ensures both fingers maintain secure contact with the object while accurately counting the number of thin, flexible objects being handled.
  • We conducted comprehensive experiments to evaluate RoTipBot, demonstrating its ability to count multiple layers of thin, flexible objects and then grasp them simultaneously in a single grasp closure.
@@ II-A Vision-based Tactile Sensors
In recent years, many vision-based tactile sensors have been developed, primarily classified into two main categories: the GelSight family [9], which processes raw images and employs photometric stereo for shape reconstruction; the TacTip family [10], which estimates contact by tracking pin displacement. Unlike TacTip, which is limited by pin resolution, GelSight offers superior resolution and detailed texture information. Our RoTip sensor belongs to the GelSight family.
The first GelSight prototype was developed in 2009 by Johnson and Adelson [11]. Following that, a set of subsequent developments were built with different focuses, such as enhanced lighting setups of GelSight-2017 [9] and GelSlim 3.0 [12], sensor reproducibility of DIGIT [13], and a compact shape for robot fingers of [14]. While previous GelSight tactile sensors capture high-resolution tactile images, their limited sensing area poses challenges for complex robotic manipulation. To address this, recent works have introduced finger-shaped tactile sensors, such as GelTip [15, 16, 17], OmniTact [18], Insight [19] and GelSight360 [20]. These sensors can be mounted on off-the-shelf parallel grippers to enhance manipulation tasks with tactile sensing.
TABLE I: Comparisons of our RoTipBot against other grippers with rotatable tactile sensors.
Feature TacRot [21] TRRG [22] RoTipBot (Ours)
Unconstrained Rotation $\times$
Omnidirectional Sensing $\times$ $\times$
Perceptible Objects In-hand Only In-hand Only Both In-hand and Out-of-hand
However, these setups possess limited degrees of freedom, resulting in a decoupling between tactile sensing and robotic dexterity. To address this limitation, several studies have explored the integration of rotational movement into tactile sensors to enhance functionality, as summarised in Table I. TouchRoller [23] was proposed to enable rapid reconstruction of large contact surfaces through surface rotation, while TacRot [21] uses a rotatable surface to facilitate in-hand object manipulation. Nonetheless, TacRot [21] is restricted to sensing only flat elastomer surfaces, and its design with the RGB strip control wire and USB camera cable bundled together limits its rotation range to ±720 degrees. This constraint makes it unsuitable for the continuous rolling motion needed to feed and grasp thin, flexible objects. The tactile-reactive roller grasper (TRRG) [22, 24] further advances this direction by integrating compliant, steerable cylindrical fingertips that allow unconstrained rotation and perception of the roller’s inner side surface. TRRG demonstrates superior performance in in-hand manipulation tasks such as cylindrical object rotation and cable tracing. However, its lack of omnidirectional sensing increases the risk of unintended collisions between the object and the non-sensing area of the grasper, particularly when interacting with objects located outside the hand. As a result, TRRG is not suitable for grasping objects with unknown locations or in scenarios where initial contact cannot be reliably guided by other sensors. In contrast, our proposed RoTip sensor integrates a tactile finger with omnidirectional sensing, enabling reliable detection of contact even under vision localisation errors. Building on this advantage, our RoTipBot employs a tactile-based grasping strategy that does not rely on precise visual alignment.
@@ II-B Thin and Flexible Object Grasping
Deformable object manipulation has been a growing research area with various applications, such as assembling cables in factories [25, 26, 27], assisted dressing and laundry folding [28, 29, 30, 31] and food production [32, 33]. These tasks mainly rely on visual perception [25, 34, 29, 33, 35]. However, the specific challenge of manipulating thin and flexible objects remains relatively unexplored. The small thickness of these objects compared to their length and width and their stacked layers complicate visual perception and grasping.
Recent approaches to manipulating thin and flexible objects focus on the control design. Several studies achieve thin object grasping via pre-grasp motions such as sliding [36] and scooping [37, 38]. In [36], thin objects are slid to the table edge and grasped from the side. In [39, 37], stable grasps of thin objects are achieved through scooping actions. However, scooping methods are generally less precise in separating and flipping exact pages or sheets, often resulting in all layers being grasped together.In [38], a prying strategy is proposed to tilt a thin object like a card to create clearance beneath it for a stable pinch grasp. Similarly, in [40], an underactuated crawler is used to tilt thin objects for pinch grasps. Zheng et al. [5] simplifies the problem by placing paper on the edge of the table for easy pickup, bypassing the challenge of grasping from the table surface.
@@
Another mechanism used for grasping delicate materials is the soft gripper. As discussed in [4], the three-axis compliance of a soft gripper provides unique benefits for grasping thin, flexible objects. In [8], a two-finger soft gripper is used to deform an object for a pinch grasp, but this approach requires knowledge of physical properties, such as friction, to model object deformation dynamics. Building on this, FlipBot introduced in [1] uses a two-finger soft gripper with a force/torque sensor to singulate and grasp thin and flexible objects by integrating exteroceptive and proprioceptive perceptions. However, it uses a force/torque sensor that only provides contact information at a single point, and can only grasp one page at a time, limiting its efficiency.
In contrast, our RoTipBot offers a tactile-based grasping strategy that ensures both fingers maintain secure contact with the object while accurately counting the number of thin, flexible objects. In this way, multiple thin, flexible objects can be grasped simultaneously in a single grasp closure. Although there are several studies [41, 42] on grasping multiple rigid objects, the problem of grasping multiple thin and flexible objects remains largely unexplored in current literature.
@@
III Hardware Setup
In this section, we introduce RoTip, a vision-based tactile sensor that can sense the entire fingertip area and rotate along the vertical axis, as shown in Fig. 2. We then present the hardware setup for handling thin and flexible objects, including the RoTip sensor.
III-A The RoTip Sensor
@@
The proposed RoTip sensor comprises three primary modules: the fixing module, the transmission module and the finger body, as shown in Fig. 2-(b), (c) and (d), respectively. The fixing module is used for secure attachment and precise positioning of the transmission module and the finger body, ensuring stability during operation. Meanwhile, the transmission module is used to drive the rotational motion, enabling dynamic functionality. Lastly, the finger body serves as the tactile interface, providing the sensor’s tactile perception capabilities. The fabrication details can be found on our website.
The specifications of the designed sensor are detailed as follows. Its payload capacity can be evaluated from two perspectives: the grasping payload, and the rolling mechanism payload. For grasping, the sensor can handle up to 3.4 kg, derived from the maximum payload of the UR5e robotic arm minus the weight of the gripper. For rolling, the RoTip sensor achieves a max payload of 20.3N when the actuator operates at a speed of 90°/s, which is sufficient for handling thin, flexible objects, such as the print paper that requires 2N of force to roll. To provide a clear visual reference, the dimensions of the RoTip sensor have been included in Fig. 2. To reduce the processing load for tactile images, the sampling frequency of the RoTip sensor has been set to 30 Hz with a resolution of 640$\times$480 pixels.
Figure 3: Hardware Setup and Coordinates. The hardware setup consists of a 6-DOF UR5e robotic arm, a Robotiq 2F-85 gripper, an Intel RealSense D435 RGB-D camera, and a laptop stand for holding the printer papers or other thin and flexible objects. The fingertips of the Robotiq gripper are replaced with two RoTip sensors. $W$ and $E$ represent the world coordinate that is located at the base of UR5e and the end-effector coordinate, respectively. $C$ and $M$ represent the camera coordinate and the marker coordinate, respectively. $T$ is the target coordinate related to the marker coordinate and used to direct the movement of the end-effector.
III-B Robotic Setup
As shown in Fig. 3, our robot setup consists of a 6-DOF UR5e robot [^1] https://www.universal-robots.com/products/ur5-robot/, a Robotiq 2F-85 adaptive gripper [^2] https://robotiq.com/products/2f85-140-adaptive-robot-gripper, and an Intel RealSense D435 RGB-D camera [^3] https://www.intelrealsense.com/depth-camera-d435/. The fingertips of the Robotiq gripper are replaced with two RoTip sensors. All the above devices are connected to a laptop equipped with an Nvidia RTX 4060 GPU. Furthermore, for evaluating the grasp in various poses, we utilise an adjustable stand mounted to the test board on the table. This stand is designed to hold the thin and flexible object intended for grasping, enabling consistent and reliable experimentation.
IV Thin and Flexible Object Grasping
In this section, we introduce RoTipBot, an approach that uses the RoTip rotatable tactile sensors to count multiple layers and then grasp them simultaneously in a single grasp closure. RoTipBot uses a two-finger gripper to handle thin and flexible objects: one passive finger holds the object, while the other active finger rotates to feed multiple sheets into the centre between the two fingers, allowing them to be grasped all at once. This method prevents pinching the paper from the same side, thereby avoiding potential damage and ensuring a more secure grasp. Though in this work we demonstrate the RoTipBot with a two-finger gripper, it is generalisable to a robotic hand with more than two fingers.
First, we use an RGB-D camera to obtain the rough grasping pose for handling thin and flexible objects. Following that, we propose a tactile-based adjustment method based on RoTip’s perception capability to adjust the pose of the end-effector, ensuring grasp stability. Finally, we propose a tactile-based paper counting approach and a continuous adjustment method to address the changes in stacked thickness during the feeding process of multiple layers of the object for grasping.
IV-A Vision-Guided Grasp Pose Estimation
In this subsection, we estimate the real-world grasping poses for squeezing and feeding thin and flexible objects with an RGB-D camera. To facilitate corner detection, we employ ArUco markers, streamlining the process and allowing us to concentrate on the primary challenges addressed by the proposed RoTip sensors: overcoming incomplete visual perception of stacked layers and enhancing dexterity in handling thin and flexible objects.
Figure 4: An overview of our RoTipBot for thin and flexible handling. (a) Vision-Based Grasping Generation: Given an RGB-D image obtained from a camera, we generate the grasp proposal for guiding the robot to contact and grasp the object. (b) Tactile-based Adjustment for Two-finger Sufficient Contact: To compensate for the noise from visual perception, RoTip’s sensing capability is then utilised to adjust the robot’s end-effector. Once both RoTip sensors are in contact with the object, the end-effector will be rotated around its $x$-axis to an angle inclined to the object for feeding and grasping. (c) Continuous Adjustment and Tactile Counting: A continuous pose adjustment approach is proposed to ensure two-finger contact while feeding multiple thin and flexible objects. Tactile sensing is also employed to count the number of fed pages, a process referred to as tactile counting. (d) Object Grasping: Finally, the gripper is closed to pick up the objects.
To simplify the detection of corners on various thin and flexible objects, we attach an ArUco marker to the corner as shown in Fig. 4-(a). The rationale for selecting the corners as the optimal locations for handling thin and flexible objects is detailed in Appendix B. With the known side length of the ArUco marker, we can estimate the transformation matrix from the camera’s coordinate system to the marker’s coordinate system, denoted as $\boldsymbol{\mathrm{T}_{C}^{M}}$. This is achieved by minimising reprojection error through non-linear optimisation. To account for estimation errors that can occur with ArUco markers in real-world applications, we use depth information to adjust the estimated $\boldsymbol{\mathrm{T}_{C}^{M}}$. First, we crop the depth map around the corners, similar to the approach in [1], to fit the object plane. Then, we adjust the rotation matrix of $\boldsymbol{\mathrm{T}_{C}^{M}}$ to be parallel to the depth-fitted plane, and we update the translation vector of $\boldsymbol{\mathrm{T}_{C}^{M}}$ based on the depth of the ArUco marker centre. Then, the target pose in the world frame represented with the transformation matrix $\boldsymbol{\mathrm{T}_{W}^{T}}$ can be calculated as:
IV-B Tactile-based Adjustment for Sufficient Contact
Given an estimated target pose from Sec. IV-A, moving directly to the target pose may result in collisions or insufficient contact due to the inherent positioning uncertainty of the vision system. Therefore, in this subsection, we introduce a tactile-based adjustment method that ensures both fingers securely contact thin, flexible objects to guarantee a secure grasp, i.e., two-finger contact. As is shown in Fig. 4-(b), the tactile-based adjustment includes two key steps, i.e., contact initialisation and pose adjustment.
IV-B1 Contact Initialisation
The end-effector is moved to a pre-contact pose, positioned with a translational offset along the $z$-axis relative to the target pose. A normal contact motion, using feedback control guides the robot towards the object, and terminates when a RoTip finger makes initial contact with the object.
We first design the pose error as $\boldsymbol{e_{\mathrm{T}}}=\ln(\boldsymbol{\mathrm{T}}^{-1}\boldsymbol{\mathrm{T}}_{\mathrm{ref}})^{\vee}$ between the observed pose $\boldsymbol{\mathrm{T}}$ to the target pose $\boldsymbol{\mathrm{T}}_{\mathrm{ref}}$. There are two steps in establishing the contact. When the gripper is far from the object, the Euclidean normal of the translation component of $\boldsymbol{e_{\mathrm{T}}}$’s is larger than a predefined threshold $\epsilon$ (set as 0.02 in our case), i.e., $\lVert\boldsymbol{e_{\mathrm{T}}}[0:3]\rVert>\epsilon$, a PD controller is used to control the robot to approach the surface of the thin and flexible object, i.e., $\boldsymbol{\mathrm{u}}(t)=\boldsymbol{\mathrm{K}}_{p}\boldsymbol{\mathrm{e}}_{\boldsymbol{\mathrm{T}}}(t)+\boldsymbol{\mathrm{K}}_{d}\,\frac{d\boldsymbol{\mathrm{e}_{\boldsymbol{\mathrm{T}}}}}{dt}(t)$, where $\boldsymbol{\mathrm{K}}_{p}$ and $\boldsymbol{\mathrm{K}}_{d}$ are $6\times 6$ diagonal gain matrices associated with the error and derivative errors at time $t$. Once the gripper is close enough to the object, i.e., $\lVert\boldsymbol{e_{\mathrm{T}}}[0:3]\rVert\leq\epsilon$, a feedforward control $\boldsymbol{\mathrm{v}}(t)$ is taken for $\boldsymbol{\mathrm{u}}(t)$ to move the end-effector towards the grasp pose to simplify the control.
Algorithm 1 Tactile-based Adjustment for Two-Finger Contact
2:Number of sufficient contact fingers: $N=0$
4:while True do
5:     if $N==0$ thenSec. IV-B1 Contact Initialisation
6:         if $\lVert\boldsymbol{e_{\mathrm{T}}}[0:3]\rVert>\epsilon$ then
8:         else
9:              $\boldsymbol{\mathrm{u}}(t)=\boldsymbol{\mathrm{v}}(t)$
10:         end if
11:     else if $N==1$ thenSec. IV-B2 Pose Adjustment
13:         Move to the refined pre-contact pose
14:         N = 0
15:     else if $N==2$ then
16:         Break
17:     end if
18:     update $\boldsymbol{e_{\mathrm{T}}}$ and $N$
19:end while